OpenStudy (anonymous):

Mrs. Grijalva hired a landscaping company to landscape her back yard. The company has a standing policy to plant 3 bushes for every tree planted. The company charges $75 for each tree planted and $20 for each bush. Mrs. Grijalva was billed for $810 after the planting was complete. How many bushes did the company plant in Mrs. Grijalva’s back yard? A. 2 B. 6 C. 18 D. 54

OpenStudy (texaschic101):

t = tree b = bushes 75t + 20b = 810 t = 3b This is not correct, because when I work this out, it does not come out right ?? I am trying to figure out another way to do this.

OpenStudy (texaschic101):

@goformit100 ....can you help

OpenStudy (texaschic101):

@moongazer ...can you look at this please

OpenStudy (texaschic101):

@UnkleRhaukus ...can you take a peek at this

OpenStudy (texaschic101):

@bahrom7893 ....help

OpenStudy (texaschic101):

@whpalmer4 ...can you help me set this up

OpenStudy (anonymous):

i ALREADY GOT IT

OpenStudy (texaschic101):

how did you set it up ?

OpenStudy (anonymous):

I didnt mines a credit recovery online assignment so i dont have to show work or setit up i just guessed and got it right

OpenStudy (texaschic101):

oh.....whatever works

OpenStudy (anonymous):

yea but thanks

OpenStudy (whpalmer4):

Okay, we can probably eliminate A as a possible answer choice immediately, because it is not a multiple of 3, and we're told they always plant 3 bushes for every tree. I suppose it is possible that on a small enough job they might not plant any trees, but that job shouldn't cost $810, either. Whatever. A is not the answer :-) We get 3x as many bushes as trees. \[B = 3T \] Bushes cost $20 each, and trees costs $75 each \[{Cost} = $20*B+ $75*T\]Total cost was $810\[810=20B+75T\] Now solve by substitution, writing \(3T\) wherever we see \(B\) in the cost equation: \[810=20(3T)+75T\] \[810=135T\]\[T=6\]Plug into formula giving # of bushes for number of trees\[B=3T = 3(6) = 18\]

OpenStudy (whpalmer4):

@texaschic101 we both made the same mistake, setting up the bush/tree relationship backward :)

OpenStudy (whpalmer4):

Another way to look at this: A tree + 3 bushes is the basic unit that they sell, and costs $75+3*$20 = $135. How many of those units do you get for $810?

OpenStudy (texaschic101):

oh.....I see now.....this problem was bugging me.....even if the asker didn't want to know how to do it, I couldn't get the problem out of my mind. Thanks for explaining it to me. I get it now.

OpenStudy (anonymous):

whats the answer

OpenStudy (whpalmer4):

@asmccash if you read my posts carefully, you'll find the answer.

OpenStudy (anonymous):

whats the answer i need to know plz